# Active robustness against the detuning-error for Rydberg quantum gates

### v1

@@ IV DER pulse under different noise sources
Figure 2: The infidelity of the DER(black-solid), TO(blue-dashed) pulses under different noise sources. From (a) to (c), for each $\epsilon_{\delta}$, the real sampling numbers are randomly distributed within the range of $[-\epsilon_{\delta},\epsilon_{\delta}]$, satisfying Gaussian, Uniform and U-shaped distributions with respective profiles shown on the top. Each point represents an average over 500 random samplings. No spontaneous decay is considered in the calculation.
@@
Accounting for the fact that the realistic detuning error induced by versatile experimental imperfections is unknown, so without loss of generality, we choose three different representations of the error distributions: Gaussian, Uniform, U-shaped, respectively for low, middle and high noise sources. The infidelity is shown as a function of detuning error $\epsilon_{\delta}$ in Fig.2 with same set of optimal pulse parameters(see Sec. III.C). By varying $\epsilon_{\delta}$, for a certain value, the realistic error number for single sampling is randomly obtained from a specific error function(top) and here we display the average result after significant samplings.
@@
Seen from Fig.2(a-c), the DER pulse always has a reduced sensitivity to the variation of $\epsilon_{\delta}$ no matter what the noise distribution is; while the TO pulse is very sensitive to $\epsilon_{\delta}$. This is expected because the DER pulse has exhibited the ability to overcome the influence from detuning deviation resulting in an improved robustness at the expense of high gate fidelities. However, we verify that, when the noise is relatively low forming as a Gaussian distribution the DER pulse has no advantages as compared to the TO pulse because it sustains a poor infidelity number above $10^{-3}$ as $\epsilon_{\delta}$ varies. While the DER pulse is a better choice when the detuning noise is large. By applying the Uniform or the U-shaped noise distributions, we clearly find the DER pulse always outperforms the TO pulse when $|\epsilon_{\delta}|/2\pi>0.5$ MHz. Therefore we remark that under a strong noise environment [64], the DER pulse protocol can outperform the TO protocol by exhibiting a more steady output with active robustness. An acceptable gate fidelity value above $0.99$ remains attainable for a large detuning error.
@@ V Potential Applications to realistic noises
In a realistic experiment, two common major error sources that severely limit the fidelity of quantum gate operations, are the finite atom temperature [65] and the laser amplitude noise [66]. The former causes significant motional dephasing of the ground-Rydberg transitions. And the latter not only leads to fluctuated Rabi frequencies, but will also affect the two-photon detuning by inducing ac Stark shifts [67]. Especially for gate schemes relying on the optimal control of pulse shapes this laser-amplitude derivation(the latter) can significantly affect the observed gate fidelity [68, 69]. Here we have developed the appealing gates with active robustness to the errors associated with the variation in two-photon detuning. That means the gates are more robust against any type of errors which has an effect on the two-photon detuning. Other technical imperfections that passively influence the gate performance are analyzed in Appendix C.
V.1 Doppler dephasing effect
So far the atomic thermal motion-induced Doppler dephasing is the most dominant error source in experiment, which ultimately limits the gate performance [70]. Atoms, whether they are warm or cold, are not stationary in the trap. The random velocity of the trapped atoms would lead to dephasing of ground-Rydberg excitation, manifested as a modification to the phase of laser Rabi frequencies $\Omega_{p}$ and $\Omega_{c}$, denoted by [71]
@@ V.1 Doppler dephasing effect
where $\epsilon_{\delta_{p(c)}}=\vec{k}_{p(c)}\cdot\vec{v}$ is the Doppler shift with $\vec{k}_{p,c}$ two participating laser wavevectors and $\vec{v}$ the atomic random velocity extracted from a Gaussian distribution with standard deviation $v_{rms}=\sqrt{k_{B}T/m}$ [72], where $T$ is the temperature and $m$ is the mass of atoms. Here we choose the laser wavelengths for the two-photon transition are, respectively, $\lambda_{p}=420$ nm and $\lambda_{c}=1013$ nm[see Fig.1(a)]. In order to diminish the Doppler dephasing, we use counterpropagating lasers instead of orthogonal ones which gives a minimal two-photon wavevector $k_{eff}=2\pi(1/\lambda_{p}-1/\lambda_{c})=8.76\times 10^{6}$ m-1. For a finite temperature $T$, this modification of laser phase equivalently means the realistic detuning fluctuation of excitation lasers felt by the atoms is a random variable which are $\epsilon_{\Delta}=\epsilon_{\delta_{p}}=k_{p}v$ and $\epsilon_{\delta}=\epsilon_{\delta_{p}}-\epsilon_{\delta_{c}}=k_{eff}v$.
Remember in Sec.IV, we verify that the DER pulse has more promising advantages when the detuning error is large, especially for Uniform and U-shaped noise distributions. As the Doppler shift follows a standard Gaussian distribution, we improve the cost function replaced by $J_{der-i}$ in the numerical optimization, towards the target of enhancing its dependence on the choice of noise type.
The first term in Eq.(19) suggests maximizing the ideal gate fidelity in the absence of any noise, same as $J_{der}$. While the second term means the average gate fidelity ¯$\bar{\mathcal{F}}(\epsilon_{\delta})$ should be simultaneously maximized, where
with variables $\epsilon_{\delta i}$ uniformly extracted from $[-\epsilon_{0},\epsilon_{0}]$. It is worth noting that the calculation of ¯$\bar{\mathcal{F}}(\epsilon_{\delta})$ is cleverly performed by introducing a set of Gaussian-type weights $\omega_{1},\omega_{2},...,\omega_{n}$ with $\sum_{i=1}^{n}\omega_{i}=1$, corresponding to a Gaussian sampling. Within an arbitrary noise range $[-\epsilon_{0},\epsilon_{0}]$, we solve a large number of discrete fidelity numbers $\mathcal{F}(\epsilon_{\delta i})$ alongside with a Gaussian-weighted random coefficient $\omega_{i}$ which finally leads to a more realistic estimation for the average fidelity ¯$\bar{\mathcal{F}}(\epsilon_{\delta})$. Such setting of $J_{der-i}$ leads to a tunable and noise-type-dependent weight $\omega_{i}$ that can be easily applied to other noise sources.
@@
Figure 3: (a) The gate infidelity of the TO(dashed-blue), DER(black), and DER-i(dash-dotted red) pulses in a wide parameter range of $T\in[0,3.0]$ mK. Inset: Fidelity versus detuning error $\epsilon_{\delta}$ for the three pulses. (b) The realistic gate fidelity versus the temperature $T$. Here, the spontaneous decays from the intermediate state $|p\rangle$ and the Rydberg state $|r\rangle$ are considered, and other imperfections are not included. The optimal gate parameters for a DER-i pulse are $(t_{1},t_{2},\omega)=(0.6508,0.9053,0.1627)$ $\mu$s. Each point represents an average over 500 random samplings.
@@
The gate infidelity caused by residual thermal motion of atoms is revealed by calculating the relationship between $\mathcal{F}(0)-\mathcal{F}(T)$ and the atomic temperature $T$, as plotted in Fig.3(a). To quantify the effect of Doppler effect, we compare the results by using the TO, DER, and DER-i pulses. We observe that the DER and DER-i pulses both benefit from a smaller dephasing error when the temperature is large outperforming the TO pulse. Inset of (a) gives a comparison of fidelity numbers under different $\epsilon_{\delta}$ values which clearly confirms that the fidelity of the DER and DER-i pulses is counterintuitively increased for larger detuning deviations. To show the gate performance in a more realistic condition, we calculate the gate fidelity by involving the effects of spontaneous decays($|p\rangle,|r\rangle$) as well as a finite temperature, as display in Fig.3(b). When $T<0.5$ mK(corresponds to $\epsilon_{\delta}/2\pi<0.3048$ MHz), the DER and DER-i pulses have no advantages compared to the TO pulse, because the TO pulse is expected to perform best when the imperfection is small. However, as the temperature increases to $T>1.0$ mK we remark that the gate fidelity using the TO pulse is dramatically decreased, and the DER pulse shows a higher tolerance there. Moreover, associated with the consideration of Gaussian-type noise the DER-i gate also outperforms the other two gates in high temperatures. For a high-fidelity gate with $\mathcal{F}$ staying over 0.99 the predicted atomic temperature can be as large as 1.5 mK(corresponds to $\epsilon_{\delta}/2\pi\approx 0.5280$ MHz) using the DER-i pulse scheme. This requirement significantly relaxes the temperature constraint for trapped atoms in optical tweezers, greatly protecting the coherence in quantum gate operation from atomic temperature effect [73].
V.2 Laser amplitude noise
Figure 4: The gate infidelity of the three pulses as a function of laser amplitude derivation $\epsilon_{\Omega}$. Here the DER-i pulse is re-optimized using a set of Uniform-type weights $\omega_{i}$(see Eq.20), resulting in the optimal gate parameters $(t_{1},t_{2},\omega)=(0.6632,0.9239,0.1658)$ $\mu$s. Note that for arbitrary $\epsilon_{\Omega}$ the realistic laser amplitude noises $\epsilon_{\Omega_{p}},\epsilon_{\Omega_{c}}$ are random, independently extracted from $[-\epsilon_{\Omega},\epsilon_{\Omega}]$ by sufficient samplings ($N$=500). Inset shows the relationship between the amplitude deviation $\epsilon_{\Omega}$ and the maximal ac Stark shifts $\epsilon_{\delta,\max}$ on the two-photon detuning. The linetype is same as in Fig.3.
@@ V.2 Laser amplitude noise
We proceed by verifying the DER and DER-i pulses are also more robust against the laser amplitude noise. Amplitude fluctuations of the excitation lasers will not only affect the Rabi frequencies $\Omega_{p}(t)$ and $\Omega_{c}(t)$, but also arise an extra two-photon detuning error $\epsilon_{\delta}$. Because, typically the laser amplitude fluctuations [74] can result in an uncertainty in Rabi frequencies as
@@
For a two-photon transition system with a large intermediate detuning $|\Delta_{0}|\gg\Omega_{p,c}$, a significant ac Stark shift is revealed to the two-photon resonance $\delta_{ac}=\frac{\Omega_{c}^{2}-\Omega_{p}^{2}}{4\Delta_{0}}$(remember $\delta_{0}=0$), which evolves with time since $\Omega_{p}(t),\Omega_{c}(t)$ are time-dependent [75]. While replacing by the fluctuated Rabi frequencies the realistic ac shift $\delta_{ac}$ takes a more complex form as
@@
where the known(first) term $\delta_{ac,0}$ can be sufficiently compensated by a fixed detuning in experiment [76]. Nevertheless, the unknown (second and third) terms induce an unknown two-photon detuning error, which is
@@
where we have ignored the higher-order terms. The maximal detuning deviation $\epsilon_{\delta,\max}$ induced by ac Stark shifts is roughly calculated by $\epsilon_{\delta,\max}=\max[\frac{\Omega_{p}^{2}+\Omega_{c}^{2}}{2\Delta_{0}}]\times\epsilon_{\Omega}$ which is linear with the increase of $\epsilon_{\Omega}$ as plotted in Fig.4(inset). For a large amplitude deviation $\epsilon_{\Omega}=0.05$, the maximal detuning error is about $2\pi\times 0.32$ MHz.
To quantify the robustness of different optimal pulses to the laser amplitude noise, Fig.4 shows the gate infidelity $\mathcal{F}(0)-\mathcal{F}(\epsilon_{\Omega})$ under different amplitude deviation rates $\epsilon_{\Omega}\in[0,0.05]$. Although the infidelity increases exponentially with $\epsilon_{\Omega}$ for all three pulses, the DER and DER-i pulses benefit from a relatively smaller infidelity number as compared to the TO pulse, confirming the active robustness towards the laser amplitude noise by our new schemes. Because the amplitude deviation rate $\epsilon_{\Omega}$ here, not only induces a laser amplitude noise that equally applies for the gate implementation, but also correlates an extra two-photon detuning error $\epsilon_{\delta}$ from the ac Stark shift effect. So for any value of $\epsilon_{\Omega}$ the robust pulses(DER, DER-i) always outperform the TO pulse for a smaller infidelity number. We also note that the DER-i pulse fails to outperform the DER pulse in this range because the induced detuning error $\epsilon_{\delta,\max}/2\pi\leq 0.32$ MHz is still very small.

### v2

@@ IV Robust pulses under different noise sources
Figure 2: The infidelity of the DER(black-solid), the DER-i(red-dotted) and the TO(blue-dashed) pulses under different types of noise sources. From left to right, for each $\epsilon_{\delta}$, the real sampling numbers are randomly distributed within the range of $[-\epsilon_{\delta},\epsilon_{\delta}]$, satisfying Gaussian, Uniform and U-shaped distributions with respective profiles schematically shown on the top. Cases of $\epsilon_{0}/2\pi=0.8$ MHz and 0.3 MHz are given in (a1-c1) and (a2-c2). Each point presents an average over 500 random samplings. No spontaneous decay is considered in the calculation.
@@
Accounting for the fact that the realistic detuning error caused by versatile experimental imperfections is essentially unknown, without loss of generality, we choose three different representations of the error distributions: Gaussian, Uniform, U-shaped, individually for low, middle and high noise sources. The infidelity is shown as a function of detuning error $\epsilon_{\delta}$ in Fig.2 with same set of optimal pulse parameters (see Table I, rows 3-7). For a certain value $\epsilon_{\delta}$, the realistic error number in single sampling is randomly obtained from a specific error function (top) and we display the average result after significant samplings.
@@
Seen from Fig.2(a1-c1), for a wide error range $\epsilon_{\delta}/2\pi\in[-0.8,0.8]$ MHz, the DER pulse always presents a steady insensitivity to the variation of $\epsilon_{\delta}$ no matter what the noise distribution is; while the TO pulse does not show any robustness and is extremely sensitive to $\epsilon_{\delta}$. This is expected because the DER pulse has exhibited the ability to overcome the influence from detuning deviation resulting in an active robustness yet at the expense of gate fidelities. Moreover, note that the DER-i pulse involving the consideration of weighted average fidelity has a much improved insensitivity as compared to the TO pulse, while also benefiting from a higher fidelity ($\mathcal{F}_{der-i}(\epsilon_{\delta}=0)\approx 0.9996$) than the DER pulse. As accounting for the noise types, we verify that, when the noise is relatively low forming as a Gaussian distribution, the DER pulse has no advantages as compared to the TO pulse because it sustains a poor infidelity above $10^{-3}$ as $\epsilon_{\delta}$ varies. The DER pulse becomes a better choice only if the detuning noise is significantly large. E.g. by applying the Uniform or the U-shaped noises, we find the DER pulse is best for $|\epsilon_{\delta}|/2\pi>0.6$ MHz. In addition, we expectedly find the improved DER-i pulse outperforms the DER pulse by bringing a more acceptable gate fidelity when $|\epsilon_{\delta}|/2\pi<0.4$ MHz, which achieves $1-\mathcal{F}_{der-i}\approx 4\times 10^{-4}$. Beyond that, although it behaviors similarly to the TO pulse, yet the DER-i pulse can show much better robustness to the detuning deviations. Therefore we remark that, under a relatively strong noise environment [68], the DER-pulse protocols can outperform the TO protocol by exhibiting an exactly steady output with perfect robustness. An acceptable gate fidelity value above $0.99$ remains attainable for a large detuning error. Moreover, the DER-i pulse is also significantly better than the TO pulse by providing both a high fidelity and good robustness.
Similar results under fine optimization are summarized in Fig.2(a2-c2). For a Gaussian-type noise the DER pulse has no advantages in fidelity because the average error felt by atoms is reduced due to the Gaussian distribution. And the DER-i pulse only outperforms the TO pulse at very large values of $\epsilon_{\delta}$. While for Uniform or U-shaped noises, we observe that, the DER pulse always preserves the most stable output when $\epsilon_{\delta}$ is varied, and the achievable infidelity has been lowered to be around $1-\mathcal{F}_{der}\approx 10^{-4}$ over the complete optimization range. In contrast, the DER-i pulse is more qualitatively similar to the TO pulse providing a larger gate fidelity. E.g. its infidelity number is close to $\sim 10^{-5}$ for small deviations. However, its active robustness to the error fluctuations is slightly weak because the infidelity exponentially increases as $\epsilon_{\delta}/2\pi\gtrsim 0.12$ MHz.
To conclude we have presented two different robust pulses DER and DER-i which provide advanced improvements in achieving higher gate fidelities and stronger robustness, especially in a noisy environment. By comparing the pulse performance we show the fine optimization with small error deviations is more favored by us due to the existence of a lower average fidelity.
@@ V Potential Applications to realistic noises
In a realistic experiment, two common major error sources that severely limit the fidelity of quantum gate operations, are the finite atom temperature [69] and the laser amplitude noise [70]. The former causes significant motional dephasing of the ground-Rydberg transitions. And the latter not only leads to the fluctuated Rabi frequencies, but will also affect the two-photon detuning by inducing ac Stark shifts [71]. Especially for gate schemes relying on the optimal control of pulse shapes this laser-amplitude derivation(the latter) can significantly affect the observed gate fidelity [72]. Here we develop the appealing gates with active robustness to the errors associated with the variation in two-photon detuning. That means the gates are in principle more robust against any type of errors affecting the two-photon detuning.
In this section, we evaluate the realistic gate performance under two major error sources. Other technical imperfections that passively influence the gate, will be analyzed in Appendix C.
V.1 Doppler dephasing error
So far the atomic thermal motion-induced Doppler dephasing is the most dominant error source in experiment, which ultimately limits the gate performance [73]. Atoms, whether they are warm or cold, are not stationary in the trap. The random velocity of trapped atoms would lead to the dephasing of ground-Rydberg excitation, manifested as a modification to the phase of laser Rabi frequencies $\Omega_{p}$ and $\Omega_{c}$, denoted by [74]
@@ V.1 Doppler dephasing error
where $\epsilon_{\delta_{p(c)}}=\vec{k}_{p(c)}\cdot\vec{v}$ is the Doppler shift with $\vec{k}_{p,c}$ two participating laser wavevectors and $\vec{v}$ the atomic random velocity extracted from a Gaussian distribution with standard deviation $v_{rms}=\sqrt{k_{B}T/m}$ [65], where $T$ is the temperature and $m$ is the mass of atoms. Here we choose the laser wavelengths for the two-photon transition of $\lambda_{p}=420$ nm and $\lambda_{c}=1013$ nm [see Fig.1(a)]. In order to diminish the Doppler dephasing, we use counterpropagating lasers instead of orthogonal ones which gives a minimal two-photon wavevector $k_{eff}=2\pi(1/\lambda_{p}-1/\lambda_{c})=8.76\times 10^{6}$ m-1. For a finite temperature $T$, this modification of laser phase equivalently means the realistic detuning fluctuation of excitation lasers felt by the atoms is a random variable which are $\epsilon_{\Delta}=\epsilon_{\delta_{p}}=k_{p}v$ and $\epsilon_{\delta}=\epsilon_{\delta_{p}}-\epsilon_{\delta_{c}}=k_{eff}v$.
@@
Figure 3: The realistic gate performance with the optimal TO(dashed-blue), DER(black-solid), and DER-i(dash-dotted red) pulses, corresponding to different optimization ranges $\epsilon_{0}/2\pi=0.8$ MHz(left column) and $0.3$ MHz(right column). (a1-a2) Real fidelity versus the detuning error $\epsilon_{\delta}$. (b1-b2) Infidelity as a function of the atomic temperature $T$. (c1-c2) The realistic gate fidelity versus $T$. Inset of (c2): The small range of $T\in[0,80]$ $\mu$K is amplified. Here, the spontaneous decays with rates $\gamma_{p}=8.4746$ MHz, $\gamma_{r}=2.7$ kHz from the intermediate state $|p\rangle$ and the Rydberg state $|r\rangle$, are considered in the calculation, and other imperfections are not included. All optimized gate parameters are given by Table 1. In (b-c), each point represents an average over 500 random samplings.
For demonstrating the improved robustness against Doppler dephasing errors, we first display the real gate fidelity for different detuning errors $\epsilon_{\delta}$, analogously to Fig.1(b1-b2) but in the presence of spontaneous decays. By varying $\epsilon_{\delta}$, if $\epsilon_{0}/2\pi=0.8$ MHz (wide optimization), it is clear that the DER and DER-i pulses can counterintuitively achieve a high fidelity $\sim 0.995$ for larger detuning deviations; however, which is at the cost of a low average gate fidelity under small deviations. Because there exists a balance between them and an improved robustness requires the cost of ideal gate fidelities. On the contrary, as for a sufficiently fine optimization, we see that both DER and DER-i pulses have dominant advantages wholly outperforming the TO pulse for any $\epsilon_{\delta}$. Because the proposed gates utilizing optimized DER or DER-i pulses spend less time in the intermediate excited states and are thus suffering from a smaller decay error than the TO pulse. In addition, by comparing DER and DER-i it is expected that the DER-i pulse shows the best fidelity $F_{der-i}\approx 0.9985$ for $|\epsilon_{\delta}|/2\pi<0.12$ MHz even outperforming the DER pulse there, because the optimization for DER-i pulse inherently has included a consideration of Gaussian-weighted noise distribution, leading to a smaller effect on the average fidelity at small $\epsilon_{\delta}$ values.
@@
The real gate infidelity caused by residual thermal motion of atoms is revealed by calculating the relationship between $\mathcal{F}(0)-\mathcal{F}(T)$ and the atomic temperature $T$, as plotted in Fig.3(b1-b2). To quantify the Doppler effect, we compare the results by using the TO, DER, and DER-i pulses for different optimization ranges. Accordingly, we choose a reasonable range for the atomic temperature variation to be $T\in[0,2.0]$ mK ($k_{eff}v_{rms}/2\pi\lesssim 0.6097$ MHz) and $T\in[0,0.6]$ mK ($k_{eff}v_{rms}/2\pi\lesssim 0.3339$ MHz) by considering the long-tail feature of Gaussian function. As $T$ increases, we observe that both the DER and DER-i pulses benefit from a smaller dephasing error for any $T$, while the DER pulse is even better because of the significantly-improved robustness against Doppler errors. Furthermore, to show the gate performance in a more realistic condition, we calculate the gate fidelity by involving the effects of spontaneous decays as well as a finite temperature, as display in Fig.3(c1) and (c2). As for a wide optimization, the DER and DER-i pulses have no advantages at low temperatures, because the TO pulse is expected to perform best when the imperfection is sufficiently small. However, as $T$ increases e.g. $T>0.5$ mK, we remark that the gate fidelity using the TO pulse is dramatically decreased, while the DER and DER-i pulses can provide higher fidelities there which strongly confirms the improved robustness against Doppler errors with them.
Finally, we calculate the real gate fidelity in the presence of Doppler errors by using fine-optimization pulses. Fig.3(c2) shows the gate fidelity of three pulses over a relatively small range of imperfections i.e. $T\in[0,0.6$] mK. In general we find the DER pulse can provide a noticeable but modest fidelity improvement with increasing temperature, while the TO pulse always performs worst. E.g. at $T=0.2$ mK the real fidelity of three pulses is $(\mathcal{F}_{der},\mathcal{F}_{der-i},\mathcal{F}_{to})=(0.9981,0.9978,0.9972)$. As taking into account of very small temperatures $T<80$ $\mu$K we see the DER-i pulse turns to perform best, e.g. $(\mathcal{F}_{der},\mathcal{F}_{der-i},\mathcal{F}_{to})=(0.9983,0.9985,0.9982)$ for $T=20$ $\mu$K. Because, the optimization for DER-i pulse has incorporated with the Gaussian-noise distribution in which the small detuning error plays the most vital role. Therefore, the optimized DER and DER-i pulses can provide a better performance than the TO pulse as long as $T$ is moderately varied within the optimization range. These results are able to relax the temperature constraint for trapped atoms in optical tweezers and greatly protect the coherence of quantum gate operation from the finite temperature effect [75].
V.2 Ac Stark shift error
Figure 4: (a-b) Estimation of the ac Stark shift error induced by the laser amplitude deviation $\epsilon_{\Omega}$ at $T=0$, corresponding to different optimization ranges $\epsilon_{0}/2\pi=0.8$ MHz and 0.3 MHz. Here the DER-i pulse is re-optimized with a set of Uniform-type weights (see Eq.17), resulting in the optimized parameters given in Table I (denoted by dark green). For each $\epsilon_{\Omega}$, the Stark shift error $\epsilon_{\delta}$ is randomly extracted from $\max[\frac{\Omega_{p}^{2}+\Omega_{c}^{2}}{2\Delta_{0}}]\times[-\epsilon_{\Omega},+\epsilon_{\Omega}]$ adding to the two-photon detuning, which affects the gate fidelity. The average result after 500 samplings is presented. The linetype is same as in Fig.3.
@@ V.2 Ac Stark shift error
In the following, we proceed by showing that the DER and DER-i pulses also have some robustness against the ac Stark shift caused by the laser amplitude deviation. Amplitude fluctuations of the excitation lasers will affect the Rabi frequencies $\Omega_{p}(t)$ and $\Omega_{c}(t)$ inducing a typical laser amplitude noise (see Appendix C). More crucially, it also correlates an extra detuning error $\epsilon_{\delta}$. Because, the laser amplitude fluctuations [76] can result in an uncertainty in Rabi frequencies as
@@
For a two-photon transition system with a large intermediate detuning $|\Delta_{0}|\gg\Omega_{p,c}$, a significant ac Stark shift will occur to the two-photon resonance $\delta_{ac}=\delta_{0}+\frac{\Omega_{c}^{2}-\Omega_{p}^{2}}{4\Delta_{0}}$(remember $\delta_{0}=0$), which evolves with time since $\Omega_{p}(t),\Omega_{c}(t)$ are both time-dependent [77]. While replacing by the fluctuated Rabi frequencies, the realistic ac shift $\delta_{ac}$ takes a more complex form as
@@
where the known (first) term $\delta_{ac,0}$ can be sufficiently compensated by a fixed detuning in experiment [78]. Nevertheless, the unknown (second and third) terms would induce an unknown two-photon detuning error, which is described by
@@
where we have ignored the higher-order terms. Here, the maximal detuning deviation $\epsilon_{\delta,\max}$ induced by ac Stark shifts can be roughly calculated by
$\epsilon_{\delta,\max}=\max[\frac{\Omega_{p}^{2}+\Omega_{c}^{2}}{2\Delta_{0}}]\times\epsilon_{\Omega}$ (22)
at $\epsilon_{\Omega}=\epsilon_{\Omega_{p}}=-\epsilon_{\Omega_{c}}$, which is linear with the increase of $\epsilon_{\Omega}$. For a large amplitude deviation $\epsilon_{\Omega}=0.05$, the maximal detuning error is about $\epsilon_{\delta}\approx 2\pi\times 0.32$ MHz.
To quantify the robustness of different optimal pulses to this extra detuning deviation $\epsilon_{\delta}$ (i.e. the ac Stark shift error), relevant results are summarized in Fig.4. We calculate the absolute infidelity $|\mathcal{F}(\epsilon_{\Omega}=0)-\mathcal{F}(\epsilon_{\Omega})|$(without the decay error) for different values of $\epsilon_{\Omega}$ at $T=0$. For each $\epsilon_{\Omega}$ the induced Stark error is randomly obtained from $\epsilon_{\delta}\in[-\epsilon_{\delta,\max},+\epsilon_{\delta,\max}]$ according to Eq.(22), serving as a two-photon detuning error for state $|r\rangle$. We treat $\epsilon_{\Omega}$ only as a quantity to measure the strength of Stark error while keeping $\Omega_{p}(t)$ and $\Omega_{c}(t)$ unvaried for avoiding the amplitude error. Since a typical amplitude deviation $\epsilon_{\Omega}$, not only induces a laser amplitude error but also correlates an extra two-detuning error $\epsilon_{\delta}$ via ac Stark shift effect. The former will be discussed in Appendix C.
Here in Fig.4(a-b), we observe that, the TO pulse does not show any robustness to the ac Stark shift error and increases exponentially with $\epsilon_{\Omega}$ as expected. While for two robust pulses DER and DER-i, the improvement of robustness is explicit. In the case of wide optimization, see Fig.4(a), we find the DER pulse is more robust to the Stark error especially for large deviations because of its inherent robustness to large detuning errors. Here the DER-i pulse is qualitatively similar to the TO pulse without having noticeable advantages. As turning to the fine-optimization case, we admit that two robust pulses are relatively worse than the TO pulse in the range of small detuning deviations because the improvement of robustness has to loss some gate infidelities. However, both of them reveal sufficient robustness against the ac Stark shift error for $\epsilon_{\Omega}>0.025$, enabling the gate infidelity to be staying below $\sim 10^{-4}$, which outperforms the TO pulse by orders of magnitude. This is achieved by the fine optimization algorithm that allows the robust pulses to have both a large gate fidelity and strong robustness.
In Sec. V.A and V.B, we have shown the active robustness of optimized DER and DER-i pulses against two realistic detuning errors arising from Doppler dephasing and ac Stark shift. As for the Doppler shift which serves as a significant two-photon detuning error, the robustly optimized pulses DER and DER-i can always outperform the non-robust TO pulse by using a fine-optimization method. While for the ac Stark shift error induced by the laser amplitude deviation, a Stark shift-robust version only exists for the larger amplitude deviations. Because two robust pulses are inherently obtained at the expense of gate fidelities at small detuning errors. While, in general, the DER and DER-i pulses can show stronger robustness to the change of amplitude deviations, e.g. for $\epsilon_{\Omega}<0.05$, the ac Stark shift error of both robust pulses can stay within $10^{-5}\sim 10^{-4}$, significantly better than the TO pulse which exponentially increases from $10^{-7}$ to $10^{-3}$.